OpenStudy (anonymous):

find a formula for the nth term {1/15, 3/20, 9/25, 27/30, 81/35}

OpenStudy (alexandervonhumboldt2):

look. the denominators of your fractions all have difference 5. the numenators have quotent 3. so your formula: \[a_n=\frac{3n-2 }{ 15 +5n }\]

OpenStudy (anonymous):

that does not work if I am correct when n=2 it will be 6-2 which is 4 on the numerator and 15+10 on the denom giving 4/25 for a2 when a2 = 3/20

OpenStudy (alexandervonhumboldt2):

oops sorry. hold on a second i will redo

OpenStudy (alexandervonhumboldt2):

a_n=3^n, where n_1=0

OpenStudy (alexandervonhumboldt2):

so the whole formula will be \[a_n=\frac{3^n}{15+5n}\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

im not sure you know what you are doing

OpenStudy (anonymous):

you realize that n=1 is the first term, n=2 is the second term, n=3 is the third term........

OpenStudy (alexandervonhumboldt2):

@freckles sequence problem